Based on this Information. Can anyone help me with the following questions. Got lost. Thank you!Estmated pop captured #that has mark recaptured year 75 1 NA 70 9 583 NA 53 6 662 3 NA 13 12 69 4 NA 17 14 61 Na 9 5 41 Ln st

14-17. By what percent is the population changing between: Years 1-2? Growing/Shrinking (circle one) by _____% Years 2-3? Growing/Shrinking (circle one) by _____% Years 3-4? Growing/Shrinking (circle one) by _____% Years 4-5? Growing/Shrinking (circle one) by _____%

18. Over the 5 year time period, what is the population growth rate, in percent?

19. Given this concerning 5 year population growth rate, you decide to investigate which of the two current threats to sifaka lemurs is responsible for this trend. You know that two of the primary threats to sifaka lemurs are tick-borne disease and deforestation. Design an observational study (or possibly a pair of studies) that you could use to understand the relative importance of these two threats.

NOTE: Be sure to specify whether you would use a control-impact, BACI, or regression approach. Also, describe how your study would deal with any potentially confounding variables.
Answer #1

10. Years 1-2

λ = 662/583 = 1.135

11. Years 2-3

λ = 69/662 = 0.104

12. Years 3-4

λ = 61/69 = 0.884

13. Years 4-5

λ = 41/61 = 0.672

As per HOMEWORKLIB RULES, I can answer up to 4 sub-parts per question.